Purity of gold is 15 carat. What is the percentage’of gold in the ornaments ?

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
24 carat purity of gold =100 ∴ 1 carat purity or gold = 100/24 15 carat purity of gold = (100*15)/24 = 62.5 %